how many liters of a 48% solution should be added to 40 liters of a 35% solution if the final mixture is to contain a 40% solution?
Answer by Boreal(15235)

x=liters of 48% solution
The trick in this is to remember the final solution is going to be 40 liters (the other part) +x added.
.48x +40(.35)=(40+x)*0.40
.48x + 14 =16 + 0.40x
0.08x =2
8x=200
x=25 liters
total solution is 65 liters
65(.40)=26 .48(25)=12 and 40 (.35)=14
25 liters of a 48% solution